# Towards Sign Distance Function based Metamaterial Design: Neural Operator Transformer for Forward Prediction and Diffusion Models for Inverse Design

### v1

@@ 6 Residual sumation (B, out, W,H)
Table S6: Training hyperparameters of the classifier-free guided diffusion model
@@
Figure S8: The MSE loss of training history of the diffsuion model
@@
The FEM simulations were executed on CPUs of the Delta machine at the National Center for Supercomputing Applications (NCSA) at the University of Illinois at Urbana-Champaign. Model training was conducted on a single NVIDIA H100 GPU on the DeltaAI machine at NCSA. Inference tasks for the forward NOT model and the inverse design model were performed on an NVIDIA A100 GPU on the Delta machine. The computational efficiency of data generation, model training, and inference is summarized in Table S7.
Table S7: Computational efficiency of data generation, model training, and inference

### v2

@@ 6 Residual sumation (B, out, W,H)
Table S6: Training hyperparameters of the classifier-free guided diffusion model.
@@
Figure S8: MSE loss of training history of the diffusion model.
@@
The FE simulations were executed on CPUs of the Delta machine at the National Center for Supercomputing Applications (NCSA) at the University of Illinois at Urbana-Champaign. Model training was conducted on a single NVIDIA H100 GPU on the DeltaAI machine at NCSA. Inference tasks for the forward NOT model and the inverse design model were performed on an NVIDIA A100 GPU on the Delta machine. The computational efficiency of data generation, model training, and inference is summarized in Table S7.
Table S7: Computational efficiency of data generation, model training, and inference.